OpenStudy (anonymous):

this 3x+2y=-5 4x-3y=16

OpenStudy (anonymous):

ok multiply the first equation by 3 straight across to get \[9x+6y=-15\]

OpenStudy (anonymous):

then multiply the second equation by 2 to get \[8x-6y=32\]

OpenStudy (anonymous):

add them and the y's will add to zero, so they drop out of the resulting equation \[9x+6y=-15\] \[8x-6y=32\] \[17x=17\] so now you know that \[x=1\]

OpenStudy (anonymous):

now that you know x is 1 you can replace x by 1 in either equation to get y. i choose the first one \[3\times 1+2y=-5\] \[3+2y=-5\] \[2y=-8\] \[y=-4\] so your answer is \[x=1,y=-4\]